# AlphaEarth Foundations: An embedding field model for accurate and efficient global mapping from sparse label data

### v1

@@ 11 Acknowledgements
We would like to thank Olaf Ronnenberger for his manuscript review and feedback, Carlos Guerra for his contributions to our dataset production infrastructure and manuscript review, Katelyn Tarrio for her assistance with the GLaNCE dataset, Maxim Neumann for his assistance preparing an earlier evaluation dataset, Sai Cheemalapati for his assistance with serving our data, Jonathan Thompson and Xiaojie Gao at Harvard Forest for testing an earlier version of our embedding fields, and the memory of The Moose for her cat support over the lifetime of this project, as best she could offer it.
@@ 13 Data and Materials Availability
We include the Cloud Score+ (Pasquarella et al., 2023) "cloud score" (cs) band as a mask with each Sentinel-2 image, binarized to 0 / 1 by thresholding at 0.5. Mask information is only used during training, no input composting or masking is performed, and mask information is not provided to the model at inference time.
@@
We use PALSAR-2 ScanSAR ("JAXA/ALOS/PALSAR-2/Level2_2/ScanSAR") imagery, which is ortho-rectified and radiometrically terrain-corrected. We include all available bands: "HH", "HV", "LIN". Most images have horizontal polarization ("HH"), vertical polarization ("HV"), local incidence angle "LIN"), and QA mask ("MSK") bands are always present, though a small subset (<8%) has only HH. We convert the intensity values from digital numbers (DN) to decibels using:
@@
To ensure that our sample also represents temporal variability in surface properties, we sample two temporal (support) periods per site. This has the added benefit of also effectively doubling the size of our training sample. The general temporal processing strategy is as follows: if there are no Wikigeo points present, a site can select any two non-overlapping periods in the sampling years. If there are points, we pick two periods to maximally allocate point date ranges that intersect those periods such that a point is allocated to only one of the two periods. Upon selection of the periods, we create a final dataset bearing all sites with all period bounds (two for each site). From this, we create our final collection of ($x$, $y$, $t_{start}$, $t_{end}$) coordinates, which resulted in a total of 10,203,798 unique rows that moved on to data source collection.
@@
For time conditional summarization following STP, we produce a temporal summary leveraging time-axial attention pooling based on a single learned query feature derived from the valid period $[t_{s},t_{e})$ following a conversion to sinusoidal timecodes. This summary is up-sampled to size L using a learned kernel. We then introduce a variational bottleneck with a key innovation: rather than collapse the output spatially, we estimate the mean direction across an $LxL$ grid of von Mises-Fisher (VMF) distributions in $S^{63}$. This bottleneck construction permits a high-degree of spatial precision without the further fine-tuning that is typical of other unsupervised embedding models, and provides a mechanism to parameterize the "smoothness" of a given embedding manifold via the VMF concentration.
AEF is trained to conditionally decode embeddings from the bottleneck. For each of $i\in M_{D}$ decoded sources, a small decoder network accepts an embedding, and a set of conditional metadata specific to that source (Figure 2B). Typically this is at least a sinusoidal timecode representing an instant in the valid period $[t_{s_{i}},t_{e_{i}})$ normalized to [0,1)[0,1)[ 0 , 1 ), though may also contain orbital geometry and metadata that is only relevant to the act of measurement, not the measurement itself. The source decoder network is applied to every location in the output grid to produce an $LxL$ reconstruction $y^{\prime}_{i}$, one for each $i\in M_{D}$ with $C_{i}$ channels. Interestingly, these decoders have the effect of generating spatially continuous predictions for an arbitrary timestamp (e.g., dense, superresolved LiDAR profiles from GEDI). We update the parameters in the entire network to minimize the error between $y^{\prime}_{i}$ and $y_{i}$, a target source frame randomly selected to intersect the valid period and potentially held out of the inputs. The error metric varies depending on the source, and accounts for spatial misregistration of the instrument, missing data, and $y^{\prime}_{i}$ vs. $y_{i}$ resolution mismatches as is the case when our nominal embedding resolution ($10m^{2}$) does not match the source’s original resolution. We minimize cross-entropy loss for categorical sources, and L1 error for non-categorical sources.
@@
To facilitate learning, in each row a frame was randomly selected from each source sequence. For sequences serving as model inputs these frames were removed from the input sequence directly. Randomly selected frames, and their corresponding metadata and timecode, are then used for computing reconstruction losses The model’s embedding output is concatenated with sensor metadata and the observation timecode, and a small decoder is applied at each pixel embedding to reconstruct the selected frame. Losses are computed against this reconstructed frame, with the nature of the loss $fi$ changing depending on the source $i\in M$ and a source-specific weight (Table S2).
@@
Reconstructions each randomly selected summarization periods $[t^{\prime}_{s_{i}},t^{\prime}_{e_{i}})$ for source $i\in M_{D}$ s.t. $t^{\prime}_{e_{i}}-t^{\prime}_{s_{i}}>4$ days. For each reconstruction objective, a different embedding corresponding to the unique summarization period is generated, and the target timestamp is normalized to this period on [0,1)[0,1)[ 0 , 1 ). We use the embedding and normalized timecode to reconstruct source i, alongside any source specific metadata specific to the act of measurement detailed in supplemental materials S15.
@@
While tuning was not performed, as expected, evaluation scores improved when batch uniformity was present, or there was no difference. For a sweep of $b\in[0,0.001,0.005,0.01,0.1]$, we found settings of $b=0$ and $b=0.1$ to be the least performant across evaluations, and $b=0.005$ to be optimal. For some evals, the difference in performance was notable e.g. GLanCE land cover max-trial linear transfer BA scores were ~66.6% for $b=0.005$ and ~64.7% for b = 0. For others, there was only a small improvement e.g. LUCAS land cover max-trial linear transfer BA scores were ~34.7% for $b=0.005$ and ~34.2% for $b=0$.
@@
If perturbation strategy (a) is used, we choose a unique, random summarization period $[t^{\prime}_{s},t^{\prime}_{e})$ s.t. $t^{\prime}_{e}-t^{\prime}_{s}>4$ days intersecting the annual period of the non-perturbed inputs. If strategy (b) is used, we choose a summarization period across the latter six months of the input sequence. If strategy ($c$) is used, we choose a summarization period across the former six months of the input sequence. We note sequence start and end times are not aligned to any calendar unit.
@@
All points and corresponding text intersecting a training row’s geometry and date range were stored with the row. During training, a random text point is selected if available, and we choose a unique random summary period $[t^{\prime\prime}_{s},t^{\prime\prime}_{e})$ s.t. $t^{\prime\prime}_{e}-t^{\prime\prime}_{s}>4$ days intersecting the annual period of the teacher model’s inputs. We condition an MLP decoder on the language model’s output with this summary period to produce an embedding aligned with the teacher model using standard CLIP loss (Radford et al., 2021).
@@
Learning utilized the Adam optimization strategy (Kingma, 2014), with a piecewise linear learning rate schedule from 0 to $1e{-4}$ over $[0,1e^{3})$ steps, then $1e^{-4}$ to 0 over [$1e^{3},1e^{5}]$. Learning hyperparameters were selected to minimize training loss while maintaining satisfactory reconstruction visual quality, stability in the contrastive and batch uniformity objectives, and desired performance on a set of diagnostic evaluations designed to assess whether embeddings could distinguish the presence of specific input sources.
@@
We used a model dimension of $DP=128$ along the precision path, $DT=512$ along the time path, and $DS=1024$ along the space path. 15 STP blocks were used in total. Implicit decoders were two-hidden-layer MLPs with a width of 512. The VMF bottlenecks utilized a fixed concentration ($\kappa$) of $8e^{3}$.
@@
We assess AEF performance using a set of evaluation datasets we derived from ten publicly-available reference datasets (Table 1). These datasets were selected to represent archetypal classification, regression, and change detection use cases, with all datasets directly linked to real-world products and applications.
@@ 2 Smallholder oil palm 200 93
Figure S13: Distribution of OpenET ensemble values. Values are reported in total mm evapotranspiration (ET) per month.
@@
We sample the 100-meter ASTER-GED product available in Earth Engine (AG100: ASTER Global Emissivity Dataset 100-meter V003; "NASA/ASTER_GED/AG100_003") (Figure S14). The original dataset was generated for the years 2000-2008. We assume surface properties are relatively stable at 100-meter resolution (though acknowledge this introduces added uncertainty in signal). We select 2017 for both the support and valid periods. We construct our sample by first arbitrarily selecting the 8.3 $\mu$m band ("emissivity_band10") as our label, and subsetting to locations for which all wavelength emissivity estimates had standard deviations < 0.05. We further subsetted to locations exclusively over land, and within $\pm$ 60° latitude. We then oversampled by selecting 330k of the remaining locations over land, and culling them with spatial proximity filtering. We then sample 2000 locations from 10 equally spaced bins between emissivity values of 0.7 and 1.0. We selected 200 train points per 0.03 sized bin, and allocated the remainder to test (Figure S15). Our final ASTER GED evaluation has a total of 2,000 training points and 15,636 test points after pre-processing.
@@
For the linear prediction approach, we follow the "RidgeClasifier" in scikit-learn (Pedregosa et al., 2011) and use a one-vs-rest approach with a pure-linear model per class. For each class in the training set, we create labels {$,1\}$ 1 } for each item in the training set, where 1-1- 1 denotes absence of the class and 111 presence. We then use ordinary least squares to fit this model and obtain the predictor. As the classes are mutually exclusive, a given example in the validation is then classified based on which of the classifiers gave the highest prediction.
@@
For a regression task, a weighted average is used to obtain the final result:
$l_{j}=\sum_{i}^{k}\frac{w_{i}*l^{t}_{ni}}{\sum_{i}^{k}w_{i}}$ (7)
where $w_{i}$ denotes the l2 distance.
@@
For unsupervised anomaly detection, we discard the train set and l2 normalize each pair of validation embeddings providing a normalized embedding before the event: , and after the event . We now take the dot product between each pair, remapped s.t. 0 = embeddings were the same, 111 = embeddings were on opposite poles:
@@
We now choose a global threshold $s$ on (0,1)(0,1)( 0 , 1 ) to binarize all $d_{j}$, and thus provide a predicted $l_{j}$ to compare to $l_{j}v$. $s$ is chosen from one of [0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9][0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9][ 0.1 , 0.2 , 0.3 , 0.4 , 0.5 , 0.6 , 0.7 , 0.8 , 0.9 ] to maximize BA over the entire validation dataset, and then all other metrics are computed using this threshold.
@@
We use precomputed CCDC features/parameters from the global Landsat-based Earth Engine collection available for 1999-2024 (“projects/CCDC/measures/v4”), which is an updated version of the Google Global Landsat-based CCDC Segments (1999-2019) dataset described in Gorelick et al. (2023). CCDC coefficients are stored as variable-length arrays and are accessible as an Earth Engine Image Collection. We select the eight harmonic coefficients in the *_coefs bands [$offset$, $t$, $cos({\omega}_{t})$, $sin({\omega}t)$, $cos(2{\omega}t)$, $sin(2{\omega}t)$, $cos(3{\omega}t)$, $sin(3{\omega}t)]$ plus the *_rmse values for seven Landsat bands (Blue, Green, Red, NIR, SWIR1, SWIR2, thermal). To match CCDC coefficients with a valid period, we choose the segment that intersects the valid period date for single-date evaluations and the middle of the specified valid period for monthly and annual evaluations. The final CCDC feature vector or “embedding” has 56 dimensions (8 coefficients for 7 Landsat bands).
@@
To maximize the performance of the MOSAIKS baseline, we tuned its hyperparameters to the evaluation set. We tried additional versions including one using a consistently annual input date range, versions that utilize composited inputs like those described in supplemental materials S19.2.1, versions that omitted all combinations of optical / radar sources, versions with [64,128,256,1024,8192][64,128,256,1024,8192][ 64 , 128 , 256 , 1024 , 8192 ] output features, and versions stacking all features through time. We found the version described here the most performant.
@@
We adapt the approach described in the Prithvi-specific example HLS Multi-temporal Crop Classification Model (Li et al., 2023), where encoder outputs are used as embeddings in a crop classification task. The multi-temporal embedding dimension is equal to 768 multiplied by the number of frames available (1 to 3) yielding 768 to 2304 dimensions.
@@
[…]
@@
[…]
@@
For regression tasks, AEF again exhibited the best overall performance in terms of both gains in $R^{2}$ and reductions in MAE (Figures S17 and S18). We find that $R^{2}$ values for AEF are always within a valid range, noting that several other approaches produced negative $R^{2}$ values i.e. worse-than-null performance for ASTER GED (emissivity prediction). OpenET ensemble (evapotranspiration prediction) was evidently more challenging where almost all other methods had negative values (clipped to -0.01 for plotting purposes; Figure S17).
@@
Though we note the small sample use cases represented here, we do note that continuous measurements, as opposed to discrete categorical labels, are often associated with field-based observational datasets, and we expect to see external comparisons including AEF on additional regression problems in the future following our data release (see Model-as-data section).
@@
We share a full set of plots detailing the performance scaling of AEF relative to other learned baselines as additional observations are added (Figure S21). Baselines were omitted when $R^{2}$ values were < 0, or for change detection evaluations when the baseline was not time-varying (SatCLIP). We note no obvious trend among the baselines, indicating methodologies that perform better on some problems more so than others. AEF generally improves monotonically as additional observations are added for 9-of-15 evaluation datasets. There is some unpredictable non-montonicity for some evals, but we do not identify any obvious grouping or regional bias.
@@
AEF’s reconstruction task relies on a noisy bottleneck to compress and extrapolate information from the sparse input sequence. Two model hyperparameters, the embedding dimension, and the channel noise parameterized by VMF $\kappa$, directly affect the capacity of this bottleneck. Lower settings of $\kappa$ in particular will also affect the smoothness of the latent (embedding) manifold due to the regularizing effect of the noise (Kingma et al., 2013). This last property is desirable when using embeddings for nearest neighbor retrieval as distances measured along a smooth (lower dimensional) manifold are more meaningful than those in a higher dimensional space (assuming the manifold hypothesis (Fefferman et al., 2016). Therefore contention exists between a smoother embedding space (lower $\kappa$), and the information capacity of the channel and therefore embedding (higher $\kappa$). We explore this in the context of embedding dimension across all methods of transfer and trial group sizes in Figure S22. The the setting used for AEF was Embedding $D=64$, $\kappa=8e^{3}$.
@@
Figure S28: Quantization results. s82 is the quantization strategy used for our released embedding fields data. The exponent on the x-axis labels indicates the quantization power.

### v2

@@ 11 Acknowledgements
We would like to thank Olaf Ronnenberger for his manuscript review and feedback, Carlos Guerra for his contributions to our dataset production infrastructure and manuscript review, Katelyn Tarrio for her assistance with the GLaNCE dataset, Maxim Neumann for his assistance preparing an earlier evaluation dataset, Sai Cheemalapati for his assistance with serving our data, Jonathan Thompson and Xiaojie Gao at Harvard Forest for testing an earlier version of our embedding fields, Eric Smit for additional proofreading, and the memory of The Moose for her cat support over the lifetime of this project, as best she could offer it.
@@ 13 Data and Materials Availability
We include the Cloud Score+ (Pasquarella et al., 2023) "cloud score" (cs) band as a mask with each Sentinel-2 image, binarized to 0 / 1 by thresholding at 0.5. Mask information is only used during training, no input compositing or masking is performed, and mask information is not provided to the model at inference time.
@@
We use PALSAR-2 ScanSAR ("JAXA/ALOS/PALSAR-2/Level2_2/ScanSAR") imagery, which is ortho-rectified and radiometrically terrain-corrected. We include all available bands: "HH", "HV", "LIN". Most images have horizontal polarization ("HH"), vertical polarization ("HV"), local incidence angle ("LIN"), and QA mask ("MSK") bands are always present, though a small subset (<8%) has only HH. We convert the intensity values from digital numbers (DN) to decibels using:
@@
To ensure that our sample also represents temporal variability in surface properties, we sample two temporal (support) periods per site. This has the added benefit of also effectively doubling the size of our training sample. The general temporal processing strategy is as follows: if there are no Wikigeo points present, a site can select any two non-overlapping periods in the sampling years. If there are points, we pick two periods to maximally allocate point date ranges that intersect those periods such that a point is allocated to only one of the two periods. Upon selection of the periods, we create a final dataset bearing all observations with all period bounds (two periods for each site). From this, we create our final collection of ($x$, $y$, $t_{start}$, $t_{end}$) coordinates, which resulted in a total of 10,203,798 unique rows that moved on to data source collection.
@@
For time conditional summarization following STP, we produce a temporal summary leveraging time-axial attention pooling based on a single learned query feature derived from the valid period $t_{s},t_{e})$ following a conversion to sinusoidal timecodes. This summary is up-sampled to size L using a learned kernel. We then introduce a variational bottleneck with a key innovation: rather than collapse the output spatially, we estimate the mean direction across an $LxL$ grid of von Mises-Fisher (VMF) distributions in $S^{63}$. This bottleneck construction permits a high degree of spatial precision without the further fine-tuning that is typical of other unsupervised embedding models, and provides a mechanism to parameterize the "smoothness" of a given embedding manifold via the VMF concentration.
AEF is trained to conditionally decode embeddings from the bottleneck. For each of $i\in M_{D}$ decoded sources, a small decoder network accepts an embedding, and a set of conditional metadata specific to that source (Figure 2B). Typically this is at least a sinusoidal timecode representing an instant in the valid period $t_{s_{i}},t_{e_{i}})$ normalized to $[0,1)$, though may also contain orbital geometry and metadata that is only relevant to the act of measurement, not the measurement itself. The source decoder network is applied to every location in the output grid to produce an $LxL$ reconstruction $y^{\prime}_{i}$, one for each $i\in M_{D}$ with $C_{i}$ channels. Interestingly, these decoders have the effect of generating spatially continuous predictions for an arbitrary timestamp (e.g., dense, superresolved LiDAR profiles from GEDI). We update the parameters in the entire network to minimize the error between $y^{\prime}_{i}$ and $y_{i}$, a target source frame randomly selected to intersect the valid period and potentially held out of the inputs. The error metric varies depending on the source, and accounts for spatial misregistration of the instrument, missing data, and $y^{\prime}_{i}$ vs. $y_{i}$ resolution mismatches as is the case when our nominal embedding resolution ($10m^{2}$) does not match the source’s original resolution. We minimize cross-entropy loss for categorical sources, and L1 error for non-categorical sources.
@@
To facilitate learning, in each row a frame was randomly selected from each source sequence. For sequences serving as model inputs these frames were removed from the input sequence directly. Randomly selected frames, and their corresponding metadata and timecode, are then used for computing reconstruction losses. The model’s embedding output is concatenated with sensor metadata and the observation timecode, and a small decoder is applied at each pixel embedding to reconstruct the selected frame. Losses are computed against this reconstructed frame, with the nature of the loss $fi$ changing depending on the source $i\in M$ and a source-specific weight (Table S2).
@@
Reconstructions each randomly selected summarization periods $t^{\prime}_{s_{i}},t^{\prime}_{e_{i}})$ for source $i\in M_{D}$ s.t. $t^{\prime}_{e_{i}}-t^{\prime}_{s_{i}}>4$ days. For each reconstruction objective, a different embedding corresponding to the unique summarization period is generated, and the target timestamp is normalized to this period on $[0,1)$. We use the embedding and normalized timecode to reconstruct source i, alongside any source specific metadata specific to the act of measurement detailed in supplemental materials S15.
@@
While tuning was not performed, as expected, evaluation scores improved when batch uniformity was present, or there was no difference. To better understand the effect of batch uniformity, we performed a sweep of $b\in[0,0.001,0.005,0.01,0.1]$. We found settings of $b=0$ and $b=0.1$ to be the least performant across evaluations, and $b=0.005$ to be optimal. For some evals, the difference in performance was notable e.g. GLanCE land cover max-trial linear transfer BA scores were ~66.6% for $b=0.005$ and ~64.7% for b = 0. For others, there was only a small improvement e.g. LUCAS land cover max-trial linear transfer BA scores were ~34.7% for $b=0.005$ and ~34.2% for $b=0$.
@@
If perturbation strategy (a) is used, we choose a unique, random summarization period $t^{\prime}_{s},t^{\prime}_{e})$ s.t. $t^{\prime}_{e}-t^{\prime}_{s}>4$ days intersecting the annual period of the non-perturbed inputs. If strategy (b) is used, we choose a summarization period across the latter six months of the input sequence. If strategy ($c$) is used, we choose a summarization period across the former six months of the input sequence. We note sequence start and end times are not aligned to any calendar unit.
@@
All points and corresponding text intersecting a training row’s geometry and date range were stored with the row. During training, a random text point is selected if available, and we choose a unique random summary period $t^{\prime\prime}_{s},t^{\prime\prime}_{e})$ s.t. $t^{\prime\prime}_{e}-t^{\prime\prime}_{s}>4$ days intersecting the annual period of the teacher model’s inputs. We condition an MLP decoder on the language model’s output with this summary period to produce an embedding aligned with the teacher model using standard CLIP loss (Radford et al., 2021).
@@
Learning utilized the Adam optimization strategy (Kingma, 2014), with a piecewise linear learning rate schedule from 0 to 1e41e{-4} over $0,1\text{e}{3})$ steps, then $1\text{e}{-4}$ to 0 over $1\text{e}{3},1\text{e}{5}]$. Learning hyperparameters were selected to minimize training loss while maintaining satisfactory reconstruction visual quality, stability in the contrastive and batch uniformity objectives, and desired performance on a set of diagnostic evaluations designed to assess whether embeddings could distinguish the presence of specific input sources.
@@
We used a model dimension of $DP=128$ along the precision path, $DT=512$ along the time path, and $DS=1024$ along the space path. 15 STP blocks were used in total. Implicit decoders were two-hidden-layer MLPs with a width of 512. The VMF bottlenecks utilized a fixed concentration ($\kappa$) of $8\text{e}{3}$.
@@
We assess AEF performance using a set of evaluation datasets we derived from ten publicly-available reference datasets representing archetypal classification, regression, and change detection use cases (Table 1).
@@ 2 Smallholder oil palm 200 93
Figure S13: Distribution of OpenET ensemble values. Values are reported in total $m$ evapotranspiration (ET) per month.
@@
[…]
@@
For linear classification, we follow the "RidgeClasifier" in scikit-learn (Pedregosa et al., 2011) and use a one-vs-rest approach with a pure-linear model per class. For each class in the training set, we create labels $\{-1,1\}$ for each item in the training set, where 1-1 denotes absence of the class and $1$ presence. We then use ordinary least squares to fit this model and obtain the predictor with $\lambda=0$. As the classes are mutually exclusive, a given example in the validation is then classified based on which of the classifiers gave the highest prediction.
For linear regression, we perform a simple least-squares fit between predictor and target variables with $\lambda=0$.
@@
For kNN regression, a simple average is used to obtain the final result:
$l_{j}=\frac{1}{k}\sum_{i}^{k}l^{t}_{ni}$ (7)
@@
For unsupervised anomaly detection, we discard the train set and l2 normalize each pair of validation embeddings providing a normalized embedding before the event: , and after the event . We now take the dot product between each pair, remapped s.t. 0 = embeddings were the same, $1$ = embeddings were on opposite poles:
@@
We now choose a global threshold $s$ on $(0,1)$ to binarize all $d_{j}$, and thus provide a predicted $l_{j}$ to compare to $l_{j}v$. $s$ is chosen from one of $[0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9]$ to maximize BA over the entire validation dataset, and then all other metrics are computed using this threshold.
@@
We use precomputed CCDC features/parameters from the global Landsat-based Earth Engine collection available for 1999-2024 (“projects/CCDC/measures/v4”), which is an updated version of the Google Global Landsat-based CCDC Segments (1999-2019) dataset described in Gorelick et al. (2023). CCDC coefficients are stored as variable-length arrays and are accessible as an Earth Engine Image Collection. We select the eight harmonic coefficients in the *_coefs bands [offsetoffset, $t$, $cos({\omega}_{t})$, $sin({\omega}t)$, $cos(2{\omega}t)$, $sin(2{\omega}t)$, $cos(3{\omega}t)$, $sin(3{\omega}t)]$ plus the *_rmse values for seven Landsat bands (Blue, Green, Red, NIR, SWIR1, SWIR2, thermal). To match CCDC coefficients with a valid period, we choose the segment that intersects the valid period date for single-date evaluations and the middle of the specified valid period for monthly and annual evaluations. The final CCDC feature vector or “embedding” has 56 dimensions (8 coefficients for 7 Landsat bands).
@@
To maximize the performance of the MOSAIKS baseline, we tuned its hyperparameters to the evaluation set. We tried additional versions including one using a consistently annual input date range, versions that utilize composited inputs like those described in supplemental materials S19.2.1, versions that omitted all combinations of optical / radar sources, versions with $[64,128,256,1024,8192]$ output features, and versions stacking all features through time. We found the version described here the most performant.
@@
We use the Prithvi 1.0 model and adapt the approach described in the Prithvi-specific example HLS Multi-temporal Crop Classification Model (Li et al., 2023), where encoder outputs are used as embeddings in a crop classification task. The multi-temporal embedding dimension is equal to 768 multiplied by the number of frames available (1 to 3) yielding 768 to 2304 dimensions.
@@
In our extreme low-shot trials, overall performance of methods was close to random chance in many cases. For 500-fold 10-shot trials, AEF error reductions were only > 1.0x in the ~90% confidence interval for 8/15 evaluations with an average range of variation $\pm 0.38\text{x}$, and for 1000-fold 1-shot trials, AEF error reductions were only $1.0\text{x}$ in the ~90% confidence interval in 5/15 evaluations with average variation $\pm 0.49\text{x}$. While the mean gain was in AEF’s favor for both settings, we consider the extreme degree of variability indicative that adequate general 1-shot or 10-shot performance remains an unsolved research frontier.
@@
Figure S16: Classification results reported in terms of balanced accuracy. Black dotted line indicates expected accuracy given random chance / number of classes. Error bars indicate $1\sigma$ ($68.27\%$) confidence interval by bootstrapping or bootstrapping and k-folds for small datasets, e.g., ethiopia_crops, canada_crops_*.
@@
For regression tasks, AEF again exhibited the best overall performance in terms of both gains in $R^{2}$ and reductions in MAE (Figures S17 and S18). We find that $R^{2}$ values for AEF are always within a valid range, noting that several other approaches produced negative $R^{2}$ values i.e. worse-than-null performance for ASTER GED (emissivity prediction). OpenET ensemble (evapotranspiration prediction) was evidently more challenging where almost all other methods had negative values (clipped to 0.01-0.01 for plotting purposes; Figure S17).
@@
Though we note the small sample use cases represented here, we do note that continuous measurements, as opposed to discrete categorical labels, are often associated with field-based observational datasets, and we expect to see external comparisons including AEF on additional regression problems in the future following our data release (see Global embeddings dataset section).
@@
We share a full set of plots detailing the performance scaling of AEF relative to other learned baselines as additional observations are added (Figure S21). Baselines were omitted when $R^{2}$ values were < 0, or for change detection evaluations when the baseline was not time-varying (SatCLIP). We note no obvious trend among the baselines, indicating methodologies that perform better on some problems more so than others. AEF generally improves monotonically as additional observations are added for 9-of-15 evaluation datasets. There is some unpredictable non-monotonicity for some evals, but we do not identify any obvious grouping or regional bias.
@@
AEF’s reconstruction task relies on a noisy bottleneck to compress and extrapolate information from the sparse input sequence. Two model hyperparameters, the embedding dimension, and the channel noise parameterized by VMF $\kappa$, directly affect the capacity of this bottleneck. Lower settings of $\kappa$ in particular will also affect the smoothness of the latent (embedding) manifold due to the regularizing effect of the noise (Kingma et al., 2013). This last property is desirable when using embeddings for nearest neighbor retrieval as distances measured along a smooth (lower dimensional) manifold are more meaningful than those in a higher dimensional space (assuming the manifold hypothesis (Fefferman et al., 2016). Therefore contention exists between a smoother embedding space (lower $\kappa$), and the information capacity of the channel and therefore embedding (higher $\kappa$). We explore this in the context of embedding dimension across all methods of transfer and trial group sizes in Figure S22. The setting used for AEF was Embedding $D=64$, $\kappa=8\text{e}{3}$.
@@
Figure S28: Quantization results. $\text{s}8^{2}$ is the quantization strategy used for our released embedding fields data. The exponent on the x-axis labels indicates the quantization power.
@@
S22.3 Production model (AEF v2.1)
The results presented in this study reflect the performance of v2.0 of the AEF model. Based on feedback we received on annual embedding fields generated with the v2.0 model, we made a number of fixes and improvements. Specifically:
  1. 1.
    We removed requirements that specific targets or input sensors be present in our training sample and re-generated our training dataset. This lead to the addition of a large number of samples from Antarctica that had previously been dropped due to limited coverage, and increased the count of our training video sequences from 8,412,511 to 10,182,450 sequences.
  2. 2.
    Independent tests revealed a performance regression for crop classification in the conterminous United States. We determined that this was related to the inclusion of NLCD in the training mixture. We addressed this by adding additional data from the USDA Cropland Data Layers (CDL) (USDA NASS, 2024) through 2023 (omitting the data from 2024) as a target, and the loss weight for NLCD and CDL was lowered from $0.50$ to $0.25$.
  3. 3.
    We identified and fixed a bug in our processing software that where incorrect handling of time codes for Sentinel-2 images acquired on January 1 resulted in divergent embedding values and visible swath artifacts in affected years.
  4. 4.
    We identified and fixed a bug in our processing software where frame sub-sampling, like that during training, was applied at inference time. Embeddings generated with AEF v2.1 now use a full year of imagery for inference.
  5. 5.
    We took further steps to mitigate tiling artifacts by applying frame-dropout to the teacher model mirroring that applied to the student.
  6. 6.
    We also achieved a reduction in subtle artifacts from multi-resolution pixel targets by modifying re-gridding to include random shifts within the grid size prior to downsampling.